OpenStudy (anonymous):

Business Math Word Problem: Out of 9,000 college students surveyed, 540 responded that they do not eat breakfast. What percent of the students do not eat breakfast?

OpenStudy (anonymous):

Okay 9000 = 100% of the students right?

OpenStudy (anonymous):

and we want to know the % of 540 out of 9000

OpenStudy (anonymous):

540 out of 9000. . . . . sounds like division, no?

OpenStudy (anonymous):

do you understand chair46?

OpenStudy (anonymous):

will i divide 540 into 9000?

OpenStudy (anonymous):

yes and we'll get .06

OpenStudy (anonymous):

and finally we convert into percent form by multiplying it by 100

OpenStudy (anonymous):

so we'll get 6%

OpenStudy (anonymous):

understand?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

np!

OpenStudy (mathstudent55):

You can use a proportion to solve this: 9000 is to 100% as 540 is to x% \(\dfrac{9000}{100} = \dfrac{540}{x} \) 9000x = 540*100 9000x = 54000 x = 6 Answer: 6%